By how much is the sum of 7/8 and 1/9 greater than 5/6?
Answer by Boreal(15235)

Common denominators of 8.9,6 are 72
so (7/8)+(1/9)=(71/72)
5/6=60/72
It is greater by 11/72, since I can subtract directly with the same denominator in each term.